OpenStudy (anonymous):

Heat Energy & Law of Thermodynamics Question: A 0.200-kg piece of lead at 100.0°C is dropped into a calorimeter. The calorimeter is a copper can of mass 5.00  10–2 kg and already contains 0.100 kg of water at 20.0°C. Calculate the final temperature if the specific heats of water, lead, and copper are 4.20  103 J/kg K, 1.30  102 J/kg K, and 4.00  102 J/kg K respectively. If someone could help work me through it, that would be great :)

OpenStudy (anonymous):

Two words Conservation of energy that is the principle of calorimeter!

OpenStudy (anonymous):

What formulas and how do I use them to solve for this?

OpenStudy (anonymous):

you think .. lets say the final temperature is t can u tell me whether t will be more than 100 or less than 100 more than 20 or less than 20?

OpenStudy (anonymous):

Less than 100 and more than 20

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

I just don't know where to start with it or how to go about solving it

OpenStudy (anonymous):

My teacher told us that we were to find the Q for the calorimeter and the water add them then solve for the T sub F using the lead's specific heat and mass.

OpenStudy (anonymous):

But that didn't give me the answer he had on the board

OpenStudy (wolfe8):

If you show what you worked and obtained, people can point out what is the problem.

OpenStudy (raffle_snaffle):



OpenStudy (anonymous):

Ok, I did this: Q=(.1)(20)(4.2x10^3)=8400 Q=(5.0x10^-2)(20)(4x10^2)=400 8400+400= 8800 8800=(T)(.2)(1.3x10^2) 338.46=T

OpenStudy (anonymous):

I might've made a small calculation error, but does the core of it seem right? I could try fiddling with the numbers. I accidentally added 8400 and 400 and got 84400 (stupid, I know). I also didn't write down the correct answers for the problems, so I'm not sure if I have the right answer now.

OpenStudy (anonymous):

I'm fairly certain the answer is something like 2.22x10^4 Also, Are you guys still there?

OpenStudy (raffle_snaffle):

I am still here. Error could possibly be a rounding error, but that doesn't matter too much

OpenStudy (anonymous):

Does the work seem right though? It's 1 in the morning and I want to go to bed.

OpenStudy (anonymous):

Yeah, decided I don't care enough for this, I have a B in the class gd'nuff. Thanks for sitting on my question and yet saying nothing. :/

OpenStudy (anonymous):

Not to you, raffle, to the other two

OpenStudy (raffle_snaffle):

np=)